aljohn is ten years older than eden. four years hence he will be twice the age of eden. how old are both of them now?
Answer by stanbon(75887)

aljohn is ten years older than eden. four years hence he will be twice the age of eden. how old are both of them now?
----
Equations:
a = e + 10
(a+4)= 2(e+4)
-------------------
Substitute for "a" and solve for "e":
(e+10+4) = 2(e+4)
e+14 = 2e+8
e = 6 (eden's age now)
----
Since a = e+10, a = 16 (aljon's age now)
========================
Cheers,
Stan H.
==============